OpenStudy (anonymous):

what is the base salary for the Bit Labs? Hours of training Monthly salary 10 1250 20 1400 30 1550 40 1700 50 1850 60 2000 70 2150

OpenStudy (anonymous):

wait, let me try work it out first.

OpenStudy (anonymous):

Yes please , im really confused in this problem :$

OpenStudy (anonymous):

yes, I got the right answer! I'm so happy lol

OpenStudy (anonymous):

now let me explain.

OpenStudy (anonymous):

Yeeeey c: okey explain

OpenStudy (anonymous):

so we know that the base salary means that 1) it's a guaranteed amount that a person would be paid no matter the position they are in. 2) on top of the base salary is the "extra amount" they will be paid for the hours they work. so lets make the base salary equal to x. 1) x + 10y = 1250 (base salary+ 10 hours of working = $1250 total we receive) 2) x + 20y = 1400 (base salary+ 20 hours of working = $1400 total we receive) Solve for y, and tell me what you get, I want you to try this for yourself. Once you obtain your y-value substitute it back into one of the beginning equations to find your x value.

OpenStudy (anonymous):

Your x value will be your base salary amount and if you have any questions, please don't hesitate to ask.

OpenStudy (anonymous):

OKay let me try and solve it

OpenStudy (anonymous):

I dont understand how can i get the y-value?

OpenStudy (anonymous):

actually, get rid of your y-value because we can directly find x.

OpenStudy (anonymous):

have you ever solved simultaneous equations?

OpenStudy (anonymous):

x+10(125)=1250 x+1250=1250 1250/1250 x=1 ? no i have not solved simulataneouse equations

OpenStudy (anonymous):

no, that's not possible for x=1 in a realistic point of view and sorry, what grade are you in if you don't mind me asking?

OpenStudy (anonymous):

You should look at this for solving simultaneous equations. http://www.teacherschoice.com.au/maths_library/algebra/alg_10.htm then you should be able to solve it, tell me what you get for your x-value afterwards.